### v1

# Step-by-Step Reasoning to Solve Grid Puzzles: Where do LLMs Falter?

Abstract
Solving grid puzzles involves a significant amount of logical reasoning. Hence, it is a good domain to evaluate the reasoning capability of a model which can then guide us to improve the reasoning ability of models. However, most existing works evaluate only the final predicted answer of a puzzle, without delving into an in-depth analysis of the LLMs’ reasoning chains (such as where they falter) or providing any finer metrics to evaluate them. Since LLMs may rely on simple heuristics or artifacts to predict the final answer, it is crucial to evaluate the generated reasoning chain beyond overall correctness measures, for accurately evaluating the reasoning abilities of LLMs. To this end, we first develop GridPuzzle, an evaluation dataset comprising 274 grid-based puzzles with different complexities. Second, we propose a new error taxonomy derived from manual analysis of reasoning chains from LLMs including GPT-4, Claude-3, Gemini, Mistral, and Llama-2. Then, we develop an LLM-based framework for large-scale subjective evaluation (i.e., identifying errors) and an objective metric, PuzzleEval, to evaluate the correctness of reasoning chains. Evaluating reasoning chains from LLMs leads to several interesting findings. We further show that existing prompting methods used for enhancing models’ reasoning abilities do not improve performance on GridPuzzle. This highlights the importance of understanding fine-grained errors and presents a challenge for future research to enhance LLMs’ puzzle-solving abilities by developing methods that address these errors [^1] Data and source code are available at https://github.com/Mihir3009/GridPuzzle.
**footnotetext: Equal Contribution
Step-by-Step Reasoning to Solve Grid Puzzles: Where do LLMs Falter?
Nemika Tyagi1∗  Mihir Parmar1∗  Mohith Kulkarni1  Aswin RRV1  Nisarg Patel1 Mutsumi Nakamura1Arindam Mitra2Chitta Baral1 1Arizona State University  2Microsoft Research {ntyagi8, mparmar3, chitta}@asu.edu
1 Introduction
Recent advancements in LLMs such as GPT-4, Gemini, Claude-3 Anthropic (2024), Llama-2 Touvron et al. (2023), and Mistral Jiang et al. (2023) have achieved remarkable performance on a wide range of Natural Language Understanding (NLU) tasks previously thought to be exclusive to human. Beyond NLU, exploring LLMs’ logical reasoning abilities Liu et al. (2021); Saparov and He (2022); Parmar et al. (2024); Patel et al. (2024) on complex tasks such as puzzle-solving is under-explored. Past attempts have been made to evaluate models on logic-intensive grid-based puzzle-solving. However, they either do not focus on evaluating LLMs Mitra and Baral (2015); Jabrayilzade and Tekir (2020) or do not evaluate LLMs independently, but rather use neuro-symbolic approaches Ishay et al. (2023) that use external specialized solvers on LLM outputs. Here, we aim to evaluate the puzzle-solving abilities of LLMs by themselves, without the use of any external logic solvers.
To understand the reasoning capabilities of LLMs, it is important to evaluate reasoning chains, rather than the final predicted answer. There have been works that evaluate reasoning chains using objective metrics such as ROSCOE Golovneva et al. (2022), CTC Deng et al. (2021), and BARTScore Yuan et al. (2021), however, they do not focus specifically on evaluating reasoning. Some prior works propose metrics for specific reasoning tasks, such as FOLIO Han et al. (2022) and ProntoQA Saparov and He (2022). However, these methods rely on reference-based evaluation, do not focus on puzzle-solving, and do not aim to identify fine-grained errors in reasoning chains. To address these limitations, we propose a reference-free manual and automated subjective evaluation of reasoning chains to understand various fine-grained errors in reasoning chains for grid-based puzzle-solving.
Figure 1: Schematic representation of proposed pipeline. Begins with the data collection of GridPuzzle dataset (top left) and evaluating various LLMs in zero-shot CoT setting (bottom left), then analyzing reasoning chains of LLMs manually to find various error types (top right) and automate this analysis process using LLM to check the correctness of reasoning chain by finding errors (bottom right).
Motivated by Mitra and Baral (2015), we first develop GridPuzzle (Figure 1), a comprehensive evaluation dataset consisting of grid-based puzzles with grid-size of $3\times 4$, $3\times 5$, $4\times 4$, $4\times 5$, and $4\times 6$ with three levels of difficulty (easy, medium, and hard). Then, we evaluate LLMs including GPT-4, Gemini-Pro, Claude-3, Llama-2, and Mistral on GridPuzzle in zero-shot-CoT setting (Figure 1). Experimental results show that LLMs do not fare well and achieve a maximum of $5.1\%$ accuracy.
To investigate reasoning chains, we manually analyze them (Figure 1) to find fine-grained errors (further details in section 3.3). Based on this, we propose a new error taxonomy comprising five broad categories, and nine fine-grained sub-categories (Tables 1 and 2), providing deeper insights into the primary causes of the LLMs’ reasoning failures. However, scaling manual analysis to a larger set is time-consuming and laborious. Hence, we propose to leverage LLMs as auto-evaluators by creating prompts that utilize error taxonomy to automate the analysis of reasoning chains and help in identifying errors (Figure 1). While evaluating w.r.t. manual annotation, our auto-evaluator model achieves $\sim 86\%$ agreement, hence providing quality error categorization.
Beyond identifying errors and the accuracy of the final answer, we propose PuzzleEval, an LLM-based framework to evaluate reasoning chains for grid-based puzzles. PuzzleEval involves a multi-stage evaluation using GPT-4o. First, we identify key logical conclusions from the reasoning chain; second, we extract key logical concepts from these conclusions; and finally, we measure the presence of these logical concepts in the final gold answer to assess the correctness of the reasoning chain. Evaluating reasoning chains based on error categorization and PuzzleEval reveals interesting findings such as LLMs show lower accuracy despite having more error-free reasoning steps, open-source models lack reasoning skills compared to closed-source models, and the most dominant error categories are wrong reasoning and elimination. Additionally, we employ existing prompting methods such as Plan-and-Solve and Self-discover, demonstrating that these methods do not improve performance on GridPuzzle. We believe that our findings will inspire future work in the automated in-depth evaluation of reasoning chains for broader reasoning tasks and enhance the reasoning abilities of models.
2 Related Work
Puzzle-solving Task
Puzzle-solving task provides detailed insights into LLMs’ logical reasoning. Giadikiaroglou et al. (2024) categorize puzzles into (1) rule-based and (2) rule-less puzzles. Rule-less puzzles include riddles Lin et al. (2021), MCQs Zhao and Anderson (2023), programming puzzles Schuster et al. (2021), and commonsense reasoning puzzles Gu et al. (2023); however, in our work we focus on rule-based puzzles. In rule-based puzzles, past attempts have explored Sudoku Noever and Burdick (2021), Rubik’s Cube, 8-puzzle, Game of 24 Yao et al. (2024), crosswords Yao et al. (2024), chess puzzles Feng et al. (2024), card games Gupta (2023), and BoardgameQA Kazemi et al. (2024). However, grid-based puzzle solving is under-explored. Mitra and Baral (2015) proposed a grid-based puzzle dataset, but this work does not provide any insights into the performance of recent LLMs. Motivated by this, we propose a systematically curated grid-based puzzle dataset, GridPuzzle, and also provide a detailed evaluation of various LLMs in puzzle-solving.
Automatic Evaluation of Reasoning Chains
Previous works Dalvi et al. (2021); Saparov and He (2022); Han et al. (2022) have focused on reference-free evaluation, which is not reliant on gold-reasoning chains. Recently, ROSCOE Golovneva et al. (2022) proposed a suite of metrics to measure the semantic consistency, logicality, informativeness, fluency, and factuality of reasoning chains, while the ReCEval framework Prasad et al. (2023) evaluates reasoning chains based on two key properties: correctness and informativeness. Recent evaluation methods such as LLM evaluation Chiang and Lee (2023) and G-Eval Liu et al. (2023) leverage LLMs to measure the quality of reasoning chains. LLM evaluation involves presenting task instructions and a text sample to LLMs, asking them to rate the sample’s quality on a 5-point Likert scale, whereas the latter incorporates automatic chain-of-thought generated by the LLM describing the detailed evaluation steps. Additionally, Tyen et al. (2023)’s attempt to use GPT-4 as an evaluator in a few-shot setting, shows that evaluating reasoning chains remains a challenge. Furthermore, AutoRace (Automatic Reasoning Chain Evaluation) Hao et al. (2024) proposed a fully automated approach for evaluating reasoning chains that adapt to different tasks without human effort. However, these methods do not evaluate reasoning chains at the level of fine-grained error types and do not provide detailed task-specific insights. To address this, we propose LLM-based reference-free evaluation methods that identify fine-grained errors and assess the correctness of generated reasoning chains.
3 Evaluation of Reasoning Chains
3.1 GridPuzzle
To develop this dataset, we extract logic grid puzzles of various grid sizes from Puzzle Baron’s Logic Puzzles [^2] https://logic.puzzlebaron.com/. Specifically, we compile logic grid puzzles of size $3\times 4$, $3\times 5$, $4\times 4$, $4\times 5$, and $4\times 6$. Each grid size has three levels of difficulty (easy, medium, and hard) except $4\times 6$. This particular grid size has only two difficulty levels (Easy and Medium). Statistics corresponding to each grid size are presented in Figure 1 (top left).
Error
Category
Description
WW Wrong Premise and Wrong Conclusion
WR Wrong Premise and Right Conclusion
RW Right Premise and Wrong Conclusion
RR Right Premise and Right Conclusion
NC No Conclusion statement or no reasoning involved
Table 1: Proposed error taxonomy for broad categories based on manual analysis. If a statement starts with “so, therefore, hence, this means, this implies, etc.” and/or is not followed by any premise, consider the previous statement’s conclusion or the previous NC as the premise.
Category Source Sub-Category Description
Wrong Premise or No Conclusion From the clues (Example: From clue 4,….) (1) Hallucination When information is completely out of context and not present in clues.
(2) Incomplete Information Lacks necessary information to make a particular conclusion.
(3) Assumptions Statements not derived from clues directly; might include assumed information relevant to the clue.
Derived Conclusions using clues given in puzzle which was not inherently given in the clues. (4) Error Propagation Premise derived from a previous incorrect conclusion.
(5) Incomplete Information Lacks necessary information to make a particular conclusion.
(6) Wrong Assumption The derived assumption is incorrect.
Wrong Conclusion Derived using the premise (which itself is either taken directly from the clues or derived) (a) Wrong Reasoning The reasoning is incorrect, regardless of the premise’s accuracy.
(b) Error propagation Conclusion is incorrect due to an erroneous premise.
(c) Wrong Elimination All premises are present, but not all conclusions are correctly derived.
Table 2: Proposed error taxonomy for sub-categories based on manual analysis. These sub-categories are defined for cases where either the conclusion or premise is incorrect (“RW” or “WR”) or both are incorrect (“WW”). For “WW”, the error sub-categories might appear in any combinations between (1-6) and (a-c) such as ‘1a’, ‘4b’, or ‘6c’.
3.2 Manual Evaluation
To explore where exactly these LLMs falter in performing reasoning, we conduct a detailed manual analysis of the reasoning chains generated by them while solving grid-based puzzles. Our manual analysis process consists of three steps. First, we begin by segmenting the reasoning chains into individual sentences, allowing us to categorize errors more precisely. Second, we identify the premise and conclusion for each sentence and determine their respective correctness. We refrain from subdividing sentences into multiple premises or conclusions to maintain simplicity for finding errors. At last, each sentence is categorized as either containing a single premise and conclusion or being a declarative statement without a conclusion. Then, we begin assessing potential issues or errors in the reasoning chains. Now, we follow an exhaustive approach to create fine-grained error categories. We begin with 30 reasoning chains (6 puzzles x 5 reasoning chains from LLMs) to manually identify potential errors. Next, we categorize these errors in a structured format. We then add another 30 reasoning chains to see if any new types of errors emerge. If new errors are identified, we refine our categories accordingly. This process is repeated until we evaluate a total of 150 reasoning chains and no new types of errors are found. Based on this method, we have carefully filtered and categorized several errors made by LLMs, presenting them as five broad categories and nine sub-categories.
3.3 Proposed Error Taxonomy
Broad Categories
As shown in Table 1, we present five main categories: “WW” - Wrong Premise Wrong Conclusion, “WR” - Wrong Premise Right Conclusion, “RW” - Right Premise Wrong Conclusion, “RR” - Right Premise Right Conclusion, and “NC” - No Conclusion. These acronyms of broad categories are self-explanatory. For instance, the category “WW” comprises sentences where the sentence consists of a wrong premise as well as a wrong conclusion. Interestingly, we also find the “WR” category consists of instances where a wrong premise still leads to a correct conclusion. Additionally, sentences containing only information from clues or premises from previous steps fall under “NC”. We conduct further investigation as to why the premises and conclusions become incorrect.
Sub-categories: Wrong Premise
As shown in Table 2, we identified the source of the premise to determine the origin of errors: (i) ‘From Clues’ – where the premise is directly borrowed from one of the clues without any further reasoning, and (ii) ‘Derived’ – where the premise is inferred from either the clues or the previous conclusions. From Table 2, there are six possible reasons associated with two different sources for the wrong premise. When the premise originates from the source (i), we find three types of errors: Hallucination – When some factual information from the clues is distorted or completely made up; Incomplete information – When the information is correctly borrowed from the clues but it is not sufficient to make a particular conclusion; Assumptions – This is a special category where the premise is not derived but also not given exactly in the clues. It is often related to one of the clues and is of the form, “Let’s assume” or “Assuming that.” When source is derived, we find three different errors: Error Propagation – This occurs when a previously incorrect conclusion becomes the basis for a flawed premise, thereby extending the error from one conclusion to the next; Incomplete information – When the derived premise is not sufficient to make a particular conclusion; and lastly, Wrong Assumption – When the LLM reasoner clearly states that a premise was an assumption but it was incorrectly derived.
Sub-categories: Wrong Conclusion
As shown in Table 2 (source), conclusions are always logically derived from a fixed set of premises. For having a wrong conclusion in any reasoning step, we find 3 errors responsible: Error Propagation – When a conclusion is wrong strictly due to some error in the preceding premise; Wrong Elimination – When the conclusion is wrong because the LLM reasoner failed to eliminate all the unfit choices correctly. This case is specific to the grid-based puzzle task but is inherently an erroneous deduction on the LLM’s end; Wrong reasoning – The remaining incorrect conclusions that did not fit in the above categories are classified under this label.
Examples of reasoning chain evaluated by GPT-4o
Sentence:"Therefore, Zeno must be 69%, and UCLA must be 62%.",
Premise: "If Zeno were 55%, there would be no score 7% lower than 55% for UCLA.",
Conclusion: "Zeno must be 69%, and UCLA must be 62%",
Explanation: "The conclusion is incorrect as UCLA is already known to be 62% from clue 2.",
Error Category: "RW.",
Premise: "A",
Sentence:"Since the third performer used flashlights, it must be either Lora or Carmen.",
Premise: "The performer who used flashlights was either Lora or Carmen.",
Conclusion: "The third performer must be either Lora or Carmen.",
Explanation: "The conclusion is based on the incorrect premise that the third performer used flashlights.",
Error Category: "WW",
Sub Category "4B",
Table 3: Examples of reasoning chain evaluated by Auto-evaluator (GPT-4o).
3.4 Automated Evaluation
Manual analysis of reasoning chains provides a detailed categorization of errors; however, it is tedious and, therefore, challenging to scale for the entire dataset. However, analyzing the distribution of errors from our proposed taxonomy on the whole dataset is also crucial in understanding the shortcomings of LLMs’ reasoning ability. Thus we develop an LLM-based auto-evaluator to automate the process of error evaluation. To this end, we prompt the GPT-4o model to identify and categorize errors in the given reasoning chain. Our prompt consists of a system instruction followed by a user prompt containing the reasoning chain to be evaluated along with the original puzzle and its gold solution. The system prompt can be further dissected into 3 key components: the instructions, the knowledge, and an exemplar. The instruction contains all the rules that the GPT-4o needs to follow to conduct accurate evaluation and error categorization of the reasoning chains. It incorporates similar sequential steps used during the manual evaluation of reasoning chains along with the required output format. The knowledge has a detailed description of our error taxonomy including the broad and sub-categories. We also provide a preference order for selecting categories along with the description to minimize any ambiguity in the evaluation process. Lastly, the exemplar consists of a puzzle, its correct solution, the original model-produced reasoning chain, and the manually evaluated reasoning chain with our error categories. We termed this LLM-based evaluator as “Auto-evaluator”. Appendix B provides the structure of the Auto-evaluator prompt.
Using the Auto-evaluator, we evaluated a total of 1,370 reasoning chains generated by five different LLMs for solving 274 puzzles. The application of our Auto-evaluator to this large dataset allowed us to analyze the distribution of error categories on a broader scale. To validate the accuracy of the evaluations performed by the Auto-evaluator, we randomly sampled 20 reasoning chains from the manually evaluated set. The authors then compared their error category assignments to those given by the Auto-evaluator. The agreement score for the total number of reasoning steps between the manual evaluation and the GPT-4o evaluation is $\sim 86\%$. Table 3 shows the example of reasoning steps evaluated by GPT-4o.
4 Experimental Steup
4.1 Experiments
We evaluate a range of closed-source LLMs including GPT-4-Turbo, Claude-3-Opus, and Gemini-Pro, and open-source models Llama-2-13B, and Mistral-7B-Instruct on GridPuzzle in the Zero-shot-CoT setting Kojima et al. (2022). Our GridPuzzle dataset consists of a set of instances denoted as $\mathcal{P}={<p_{n}^{i\times j},a_{n}>}$, where $p_{n}^{i\times j}$ is $n^{th}$ puzzle instance with grid size of $i\times j$ and $a_{n}$ as a gold answer. We prompt each LLM to generate a reasoning chain before predicting answer $\hat{a}$. To evaluate each model in the Zero-shot-CoT setting, we provide $<I,p_{n}^{i\times j}>$ as input to the model and predict an answer $\hat{a}$ where $I$ is a natural language instruction. The evaluation is conducted on the OpenAI, Google, and Anthropic model versions released in April 2024 with temperature setting 0 for deterministic predictions. NVIDIA A100 GPUs are used for conducting the inference of open-source models with a batch size of 4. The example prompts used for these experiments are provided in Appendix A.
Figure 2: The process of calculating PuzzleEval metrics is described above. The reasoning chains are produced by our 5 LLM models and the gold solution is taken from our GridPuzzle dataset.
4.2 Metrics
Accuracy
We use accuracy to demonstrate the capability of LLMs in solving grid-based puzzles based on their ability to predict the final answer. To calculate this metric, we use the LLM-generated final answers and compare them with the available gold solution. The predicted answers and the gold solution are in the form of tables with the number of rows and columns equal to the grid size of the puzzle. We perform an Exact Match (EM) to compare the two tables and mark them as correct only when all the entries of the tables match. See the example of the final answer table in Appendix C.
Figure 3: Performance of 5 different LLMs in terms of accuracy on the GridPuzzle dataset.
PuzzleEval
We developed this LLM-based metric to assess step-by-step reasoning chains and provide a correctness score for each step, as well as the Average Correctness Score (ACS) for the entire chain. PuzzleEval is a reference-free metric specifically designed for assessing reasoning chains generated for grid-based puzzle tasks. It evaluates the correctness of each step in the reasoning chain and reports the score using only the final answer table provided as the gold solution, without requiring any comparison to a gold reasoning chain.
As shown in Figure 2, PuzzleEval consists of a three-stage pipeline to evaluate any reasoning chain. First, we prompt GPT-4o to label all the steps sequentially to account for any discrepancies in the different formats of reasoning chains produced by various models, and to extract only the final conclusions from each step. This stage is crucial as it removes the portion of a step where the models just reiterated clues or previous conclusions. Second, we instruct the model to extract the pair-wise relation of elements from the puzzle that have been either accepted or rejected in the extracted final conclusions. If the extracted conclusion is "Sam is assigned to the year 2015 but not 2014.", these pairs are of the form “Sam – 2015” or “Sam – not 2014”. Third, we provide the gold solution table and ask the model to check if these accepted or rejected pairs match the given information. As per the validation, the pairs extracted from every step are marked as correct or incorrect. After obtaining this information for each step the correctness score is calculated by adding up all the correct and incorrect steps (correct pairs are marked 1 and incorrect pairs are marked 0) divided by the total number of pairs in each step. Finally, the ACS is determined by adding up all the correctness scores from each step and dividing by the number of steps to capture the overall quality of the reasoning chain. Hence, PuzzleEval provides ACS for each reasoning chain in range of 0 to 1.
5 Results and Analysis
5.1 Objective Evaluation
To evaluate the performance of LLMs when solving grid-based puzzles, we assess the outputs of 5 LLMs using the accuracy and PuzzleEval. As shown in Figure 3, we found that all the models have low performance on the GridPuzzle dataset in terms of accuracy. The smaller open-source LLMs completely failed at the puzzle-solving task, with LLama-2 solving only one puzzle correctly. Close-source models with significantly larger parameter sizes also exhibited poor performance. GPT-4 had the highest accuracy at only 5.11% (14 puzzles out of 274). Despite the overall low performance of all LLMs, the closed-source models perform marginally better. We evaluate the quality of the reasoning chains using PuzzleEval. Table 4 provides the ACS for each grid size available in the GridPuzzle. Surprisingly, compared to the accuracy, the performance of the models with PuzzleEval was significantly better as shown in Table 4. The ACS lie in the range of 0.26 to 0.64 across all grid sizes. This higher score can be attributed to the partial correctness of reasoning chains when solving the grid-puzzle task. The disparity between metrics shows that evaluating only final answers doesn’t fully capture LLMs’ effectiveness in complex logical tasks like grid puzzles.
Model 3 x 4 3 x 5 4 x 4 4 x 5 4 x 6 Avg
Llama 0.45 0.46 0.46 0.42 0.28 0.41
Mistral 0.29 0.26 0.27 0.26 0.27 0.27
Claude 0.60 0.56 0.52 0.55 0.46 0.54
Gemini 0.60 0.64 0.54 0.52 0.62 0.58
GPT-4 0.61 0.62 0.56 0.54 0.60 0.59
Table 4: The results for PuzzleEval on the different grid sizes available in GridPuzzle dataset in terms of ACS.
With the increase in the sizes of the grids, the complexity of the puzzles also rises, leading to a depreciating performance by the LLMs with larger grids. Overall the performance of larger LLMs was much better than the small open-source models. Mistral-7B performed the worst in PuzzleEval which is in accordance with its low accuracy score. GPT-4 and Gemini models surprisingly have similar PuzzleEval scores (0.59 and 0.58 respectively) despite their large difference in accuracy. This difference in PuzzleEval could be attributed to the relatively shorter reasoning chains (fewer reasoning steps) produced by Gemini (an average of $14.91$ steps) compared to GPT-4 (an average of $20.66$ steps). Shorter reasoning chains may reduce the number of errors that occur while solving the puzzle. It is interesting to note that the smaller LLMs have consistently low performance with the increase in the grid size of the puzzles but the larger LLMs have mixed performance.
5.2 Reasoning Chain Evaluation
Figure 4: The percentage distribution of the broad error categories across the combined reasoning steps of all 5 LLMs. The total number of steps generated by each model is provided inside the round brackets below the model names.
Figure 5: The first five sub-figures in the above section show the error Sub-category distribution over five LLMS. The last sub-figure denotes the top 10 error Sub category distribution across all model reasoning steps.
The relative distribution of the broad error categories over the collective reasoning steps for each model is given in Figure 4. It is important to note that, despite using the same zero-shot-CoT setting, the GPT-4 and Llama-2 used significantly more reasoning steps ($>5.5k$ steps) to solve the 274 puzzles compared to the other three models ($\sim 4k$ steps). The distribution of error sub-categories for each model is presented as heatmaps in the first five sub-figures in Figure 5. Here, we present several findings based on the evaluation of different error category distributions across GridPuzzle.
Majority of reasoning steps are error-free.
Figure 4 shows that most reasoning steps for each model fall into the “NC” error category, indicating that many steps reiterate the facts or clues from the initial puzzle rather than focusing on reasoning. Over 55% of Gemini-Pro’s reasoning steps fall into this category, the most among all models, suggesting that Gemini spends the fewest steps on actual reasoning. The "RR" category comprises over 46% of GPT-4’s reasoning steps, highlighting its strong reasoning ability. This higher number of correct reasoning steps correlates with GPT-4’s higher PuzzleEval score, reflecting its overall effectiveness.
Why is the accuracy low despite the reasoning chains being mostly error-free?
The disparity between accuracy and PuzzleEval arises from the relative location of errors within the reasoning chains. It has been observed that “RR” category reasoning steps mainly occur in the initial half of the chain, leading to a high overall PuzzleEval score. Conversely, errors in the “RW”, “WR”, and “WW” categories typically occur in the latter half of the chain, resulting in incorrect final answers and lower accuracy scores. Based on our error taxonomy, the “RW”, “WR”, and “WW” broad error categories have been further dissected into $6\times 3$ error sub-categories, with their distribution across reasoning steps shown in Figure 5.
Dominant error sub-categories and their correlation with the broad categories.
The most common error sub-category across all heatmaps appears to be the “-” category which refers to the absence of errors. All the reasoning steps with “NC” and “RR” classifications fall in this category. To observe the actual overall trend across all 5 LLMS, the top 10 most common error sub-categories have been listed in the last sub-figure of Figure 5. The top categories ‘a’ and ‘c’ refer to the Wrong Reasoning and the Wrong Elimination sub-categories under the “RW” category. These errors arise when the premise is correct but the LLMs fail to make accurate deductions from it. A number of the top 10 sub-error categories (‘1a’, ‘4a’, ‘4b’, ‘4c’, ‘5a’, ‘5c’, and ‘6a’) emerge from the “WW” category.
For the categories, ‘4a’, ‘4b’, and, ‘4c’ the errors in the premise are propagated from errors in previous reasoning steps showing how initially occurring errors in the chain can lead to more dependent errors. The ‘4b’ error category is the one where this behavior is maximized as here both the premise and conclusions were wrong because of previously propagated errors. The ‘5a’ and ‘5c’ errors occurred due to the incompleteness or lack of information in the premise and wrong reasoning or elimination in the corresponding conclusions. The ‘1a’ kind of error occurred when the premise consisted of hallucinated information. The only sub-category from the “WR” category making it in the top 10 is the ‘3’ category which is caused due to wrong assumptions in the premise. It can be noted here that the reasoning steps of the “WR” category do not deteriorate either of the evaluation metrics, as the conclusions ended up being correct, but rather indicate the inconsistency of the LLMs in reasoning over puzzle-solving.
Mitigation Strategy Accuracy PuzzleEval
Baseline 12 0.61
Plan-and-Solve 9 0.62
Self-correct 10 0.59
Self-discover 13 0.65
Feedback-Learning 10 0.59
Table 5: The results for accuracy and PuzzleEval using GPT-4-Turbo, with and without mitigation strategies for the 30 samples of $3\times 4$ grid-size.
Performance of Open-source vs Proprietary LLMs.
From the results of our objective and subjective metrics, it is evident that the open-source models have lower performance on the grid-puzzle-solving task than the proprietary models. The Llama-2 and Mistral models have the lowest accuracy values and their low performance on the PuzzleEval consistently degrades with the increase in the size and complexity of the grids. The Claude-3, Gemini, and GPT-4 models have higher values of accuracy but their performance across the grid sizes in the PuzzleEval is inconsistent. The disparity in the performance of both kinds of models can be attributed to the difference in their parameter sizes and the low instruction following capabilities of small open-source models.
Discussion on Mitigation Strategy
We conduct a case study on a subset of GridPuzzle, focusing on a 3x4 grid size, utilizing commonly employed prompting techniques to enhance the reasoning capabilities of LLMs. In particular, we use four strategies: (1) Plan-and-Solve Wang et al. (2023), (2) Self-correct Zhang et al. (2024), (3) Self-discover Zhou et al. (2024), and (4) Feedback-Learning. We updated the prompts corresponding to these techniques to include some of our major findings from the reasoning chain evaluations and error categorization analysis as precautionary instructions.
The first strategy Plan-and-Solve prompts the model to first generate a plan to solve the given problem and then follow those steps. The second strategy is inspired by the Self-correct method which uses a combination of self-verification and self-refine to improve reasoning. Next, we used the Self-discover technique which is a 2-step structured reasoning process. Lastly, we created our prompting technique called “Feedback-Learning” by providing specific feedback system instructions to the LLM based on our error taxonomy. The detailed prompt structure is described in Appendix D and the results of these strategies are in Table 5. It is evident from the results that prompting-based strategies are not sufficient to significantly improve the LLM reasoning on the grid-puzzle-solving task. Compared to the rest of the strategies, Self-Discover marginally improves the performance on both accuracy and PuzzleEval. These results indicate the need to develop techniques beyond prompting by having deeper insights from LLMs’ reasoning chains.
6 Conclusion
In this work, we evaluated the logical reasoning abilities of LLMs through the lens of a grid-based puzzle-solving task. We introduced GridPuzzle, an evaluation dataset of 274 puzzles with various grid sizes. From a manual evaluation of reasoning chains generated by five different LLMs on GridPuzzle, we developed a fine-grained error taxonomy with five broad categories and nine sub-categories. We then created an Auto-evaluator to automate the identification of error categories, providing broader insights into error distributions across the dataset. Additionally, we proposed PuzzleEval, a reference-free metric to objectively evaluate the correctness of reasoning chains for grid-based puzzles. Our analysis of error distributions in GridPuzzle revealed several interesting findings and insights into the logical reasoning abilities of different LLMs. We further evaluated existing reasoning-specific prompting methods, such as self-discover and self-correct, finding that they do not improve results on GridPuzzle. We believe our work offers a challenging dataset, highlights where these LLMs make mistakes, and provides insights to develop better logical reasoning systems for complex tasks such as grid puzzle-solving.
Limitations
While GridPuzzle facilitates the evaluation of LLMs’ logical reasoning abilities, the complexity of the puzzles can be enhanced by incorporating further complex grid sizes beyond 4x6. Additionally, this study can be extended to different types of puzzles, such as Sudoku, Game of 24, and commonsense puzzles. Though our study provides fine-grained error categories, it can be further refined by mapping to formal logic to identify more detailed and atomic errors, offering a deeper understanding of LLMs’ reasoning failures. Although we propose an effective automatic method for error identification to reduce manual analysis, exploring other automated methods using smaller-scale supervised learning could be a promising future research direction. We also note that this research is currently limited to the English language and can be extended to multilingual scenarios to evaluate the logical reasoning abilities of LLMs.
Ethics Statement
We have used AI assistants (Grammarly and ChatGPT) to address the grammatical errors and rephrase the sentences.
Acknowledgement
We extend our gratitude to the Research Computing (RC), and Enterprise Technology at ASU for providing computing resources, and access to the ChatGPT enterprise version for experiments. We acknowledge support by a 2023 Spring Amazon Research Award (ARA).
Appendix A GridPuzzle dataset sample Puzzle
The GridPuzzle dataset contains 274 puzzles of various grid sizes and complexity. A sample puzzle from the dataset along with the Zero-shot-CoT prompt is described in Figure 6. All the puzzles in the dataset have a similar structure with varying numbers of clues.
Figure 6: The prompt structure of a 4 x 4 grid size puzzle from GridPuzzle dataset. Every Zero-shot-CoT prompt from the dataset consists of Puzzle, its corresponding Clues, the Instruction for solving the puzzle, along with the Gold solution of the Puzzle.
Appendix B GPT-4o Auto Evaluator
To expand the reasoning chain evaluation process we prompt the GPT-4o model with a detailed system prompt. The structure of this system prompt is elaborated in Figure 7. The 3 main components of this system prompt are the Instruction - similar to the ones given to human evaluators, the Knowledge - obtained from the error taxonomy, and an Exemplar - consisting of a Puzzle, its Gold Solution, the LLM-generated Reasoning chain, and the evaluated Reasoning Chain.
Figure 7: The prompt structure of auto-evaluator. The prompt consists of instructions to evaluate reasoning chains, Knowledge of the error category descriptions, and a human-evaluated reasoning chain example. This system prompt is followed by a user prompt with the puzzle, its gold solution, and the reasoning chain to be evaluated.
Figure 8: The top left section of the figure consists of a 3x4 sample puzzle from the GridPuzzle dataset along with the Zero-shot-CoT prompt. Right below the prompt, we have the Gold solution for the corresponding puzzle. In the top right section of the figure, we have the Model-generated Reasoning chain to solve this puzzle along with the Final Answer. In this particular instance, the reasoning chain was generated by the GPT-4 model. In the bottom half of the figure, we have the manually evaluated reasoning chain, split into individual sentences. The corresponding error categories found in each step is given the bottom right column called ‘Error Categories.’
Figure 9: The top left section of the figure consists of a 3x4 sample puzzle from the GridPuzzle dataset along with the Zero-shot-CoT prompt. Right below the prompt, we have the Gold solution for the corresponding puzzle. In the top right section of the figure, we have the Model-generated Reasoning chain to solve this puzzle along with the Final Answer. In this instance, the reasoning chain was generated by the Llama2-13b model. In the bottom half of the figure, we have the GPT-4o Auto-Evaluated Reasoning chain.The auto-evaluation is done sentence-wise and the output is in a JSON-structured format consisting of 5 components: the Sentence, the Premise, the Conclusion, the Error category and the Sub-category. The corresponding error categories found in each sentence are given in the bottom right columns called ‘Error Categories’ and ‘Sub Categories.’
Appendix C Evaluation of Reasoning Chains
In order to identify the error categories from the erroneous reasoning chains we conducted manual and auto-evaluation of the reasoning chains. The process of manual evaluation has been described in figure 8 and the process of auto-evaluation using GPT-4o has been described in figure 9.
Appendix D Mitigation Strategy Prompts
We conducted a study on the 60, 3x4 puzzles present in GridPuzzle dataset to try and improve the reasoning abilities of LLMs when solving the grid-puzzle task. We used prompt-based methods, such as the Plan-and-Solve technique, which divides puzzle-solving into planning and solving steps. We also enhanced the solver with insights from our error taxonomy. The prompt structure for this technique is given in figure 10.
Figure 10: The prompt structure for the Plan-and-Solve strategy which is split into two prompts one for planning and the other for solving the puzzle.
Next, we devised our own strategy to improve LLM reasoning by using the top error categories from our findings and teaching the LLM to rectify those mistakes. This strategy termed as Feedback-learning makes use of a detailed system prompt that acts as a feedback-providing unit followed by a basic user prompt to solve the puzzle. The prompt structure for this strategy is shown in figure 11.
Figure 11: The prompt structure for the Feedback-learning strategy. The system prompts consist of instructions regarding the major errors as well as ways to rectify those errors.
The Self-correct strategy merges Self-verify and Self-refine qualities to minimize LLM reasoning errors. It starts with solving the puzzle using a Zero-shot-CoT prompt, followed by prompting the LLM to verify and refine the solution. Finally, it integrates the model’s suggestions with insights from our error taxonomy to enhance the puzzle-solving response. The prompt structure for this strategy is shown in figure 12. The Self-Discover strategy, depicted in figure 13, proved most effective in reducing LLM reasoning errors in puzzle-solving. This approach begins by having the model analyze the problem and potential errors, follows with a list of prescribed reasoning modules, prompts the LLM to select and apply the most suitable module, and concludes by using a structured prompt to solve the puzzle.
Figure 12: The prompt structure for the Self-Correct strategy is split into 3 parts. The first prompt solves the puzzle, the second prompt verifies the solution and gives suggestions to improve the solution, and the third prompt uses these suggestions along with error taxonomy-based instructions to refine the final solution.
Figure 13: The prompt structure for the Self-Discover strategy. In the first part of this prompt the model is prompted to assess the problem and select the appropriate reasoning module to solve it. Then the module is modified to give a structured plan to solve the puzzle. In the second part, the model uses this structured plan along with instructions from our error taxonomy to solve the puzzle.